OpenStudy (anonymous):

3x-y=1 y+3=5x using substitution

OpenStudy (anonymous):

rearrange the second equation to get y=5x-3 plug this in for y in the first equation, and we get 3x-(5x-3)=1 3x-5x+3=1 -2x=-2 x=1 plugging x back into, the equation y=5x-3 gives y=5(1)-3 y=5-3 y=2 thus the solution is (1,2)

OpenStudy (anonymous):

dude thanks so much means alot